a police jeep is chasing a culprit going on a motor bike.the moto bike cosses a turning at a speed of 72km/h.the jeep follows it at a speed of 90km/h, crossing the tuning 10sec late than the bike.assuming that they move in constant speed,how fa fom turning will the jeep catch up with bikeYell............

Grade:Upto college level

1 Answers

sindudharsana sindudharsana
11 years ago

v=72kmph=72*1000/3600=20ms-1(bike)

v=90*1000/3600=25ms-1(jeep)

relative velocity=25-20=5ms-1

time=10+10s=20s

distance=velocity*time=5*20=100m